# Asymptotically normal estimators in high-dimensional linear regression

### v1

@@ 2 Model set‑up and main result
For each sample size $n\in\mathbb{N}$, consider $(p_{n}+1)$-dimensional observations $\{(Y_{i,n},\bm{Z}_{i,n})\}_{i\in[n]}$ satisfying
@@
where $\bm{\theta}_{n}\in\mathbb{R}^{p_{n}}$ and $p_{n}$ may diverge with $n$. The sequence $\{(\bm{Z}_{i,n},\epsilon_{i,n})\}_{i\in[n]}$ are i.i.d. with $\mathsf{E}(\bm{Z}_{1,n})=\bm{0}$, $\mathsf{E}(\epsilon_{1,n})=0$, $\mathsf{V}(\epsilon_{1,n})=\sigma_{n}^{2}>0$, and $\bm{Z}_{1,n}$ independent of $\epsilon_{1,n}$. Let $\bm{\theta}_{0,n}$ denote the true parameter vector. Hereafter, we suppress the subscript $n$ when no confusion can arise. Write $\bm{\theta}_{0}=(\theta_{01},\ldots,\theta_{0p})^{\top}$. Define the index sets
@@
where $\theta_{\min}:=\min_{j\in T_{0}}|\theta_{0j}|$, $\gamma>0$ is a fixed constant independent of $n$ and $p$, and $\eta_{p}$ is a positive sequence depending on $p$. Let $s_{0}:=|T_{0}|$ and $s_{\gamma}:=|T_{\gamma}|$ denote the corresponding support sizes, where $s_{0}$ may diverge as $n\to\infty$. The next assumption specifies the conditions on $\gamma$, $s_{0}$, $s_{\gamma}$, and $\eta_{p}$.
@@
    $s_{\gamma}$ is fixed and independent of $n$ and $p$.
@@
Assumption 1-(ii) states that the non-zero coefficients outside $T_{\gamma}$ are precisely the weak signals, and Assumption 1-(iii) requires that their aggregate magnitude is asymptotically negligible. Under Assumption 1, $(\bm{\theta}_{0,n},0,0,\ldots)\in\mathbb{R}^{\infty}$ lies in $\ell^{2}$ uniformly in $n$, because
$\|\bm{\theta}_{0}\|_{2}^{2}=\sum_{j\in T_{\gamma}}|\theta_{0j}|^{2}+\sum_{j\in T_{0}\setminus T_{\gamma}}|\theta_{0j}|^{2}\leq\sum_{j\in T_{\gamma}}|\theta_{0j}|^{2}+\eta_{p}^{2}s_{0}\leq s_{\gamma}\theta_{\max}^{2}+o(1)=O(1)$
as $n\to\infty$.
Let $\hat{\bm{\theta}}_{n}$ be a sparse estimator for $\bm{\theta}$ such that $\mathsf{P}(\|\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{\infty}>c_{n})\to 0$ as $n\to\infty$, where the rate $c_{n}$ satisfies that $c_{n}=o(1)$ as $n\to\infty$. Given a threshold sequence $\{\tau_{n}\}_{n\in\mathbb{N}}$, define the sequence $\{\hat{T}_{n}\}_{n\in\mathbb{N}}$ of an estimator of $T_{0}$ by
@@
For example, the lasso estimator and the Dantzig selector satisfy
$\|\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{\infty}=O_{p}(n^{-1/2}(\log p)^{1/2})$
as $n\to\infty$. Accordingly, in such cases we may take $c_{n}\asymp n^{-1/2}(\log p)^{1/2}$.
@@
The following condition holds: $c_{n}+\eta_{p}<\tau_{n}<\gamma-c_{n}$.
@@
Let Assumptions 12 hold. Then, $\mathsf{P}(\hat{T}_{n}=T_{\gamma})\to 1$ as $n\to\infty$.
We define the ordinary least squares estimator $\tilde{\bm{\theta}}_{n}=(\tilde{\bm{\theta}}_{n1},\ldots,\tilde{\bm{\theta}}_{np})$ after variable selection as the solution to the estimating equation
@@
$\bm{J}_{n}:=n^{-1}\sum_{i=1}^{n}\bm{Z}_{i}\bm{Z}_{i}^{\top},\quad\bm{J}:=\,\mathsf{E}[\bm{Z}_{1}\bm{Z}_{1}^{\top}].$ (2.1)
In addition, we impose the following conditions.
@@
    There exists $\lambda_{*}>0$ and $\Lambda_{*}>0$, independent of $n$ and $p$, such that $\lambda_{*}\leq\Lambda_{\min}(\bm{J}_{T_{\gamma},T_{\gamma}})\leq\Lambda_{\max}(\bm{J}_{T_{\gamma},T_{\gamma}})\leq\Lambda_{*}$.
@@
    $\bm{Z}_{1},\ldots,\bm{Z}_{n}$ are i.i.d. copies of a sub-Gaussian random vector $\bm{Z}$.
@@
    $\epsilon_{1},\ldots,\epsilon_{n}$ are i.i.d. copies of a sub-Gaussian random variable $\epsilon$.
Under Assumption 3-(ii), there exist positive constants $K_{0}$ and $K_{p}$ such that $\|Z_{j}\|_{\psi_{2}}\leq K_{0}$ for $j\in[p]$ and $\|\bm{Z}\|_{\psi_{2}}\leq K_{p}$, where $K_{0}$ is independent of $n$ and $p$, while $K_{p}$ may depend on $p$. Under Assumption 3-(iii), there exists a positive constant $K_{\epsilon}$ such that $\|\epsilon\|_{\psi_{2}}\leq K_{\epsilon}$, where $K_{\epsilon}>0$ is independent of $n$, $p$, and $s_{0}$. Hereafter, we use these constants $K_{0}$, $K_{p}$, and $K_{\epsilon}$.
Remark 3.
Under Assumption 3-(ii), there exists a universal constant $C>0$ such that $\|\bm{Z}_{S}\|_{\psi_{2}}\leq CK_{0}|S|^{1/2}$ for any $S\subset[p]$ whose cardinality is fixed and independent of $n$ and $p$.
@@
Let Assumptions 1 and 3 hold. Then, $\|\bm{J}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{T_{\gamma},T_{\gamma}}\|_{\mathrm{op}}=o_{p}(1)$ as $n\to\infty$.
@@
Fujimori and Tsukuda (2026) analyze asymptotic normality for high‑dimensional estimators via weak convergence in $\ell^{2}$, including settings with a possibly infinite‑dimensional nuisance parameter and dependent observations. Their analysis is confined to the fixed-sparsity regime, which materially streamlines the analysis. In contrast, we allow the support size $s_{0}\to\infty$ while keeping the number $s_{\gamma}$ of coefficients of moderate magnitude fixed.
To prove Theorem 2.3, it suffices to establish Lemma 2.4; see, for example, Chapter 1.8 of van der Vaart and Wellner (2023).
Lemma 2.4.
Let Assumptions 13 hold.
  1. (i)
    For every $\eta>0$, $\mathsf{P}(\|n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}^{c}}\|_{2}^{2}>\eta)\to 0$ as $n\to\infty$.
  2. (ii)
    For every $\bm{u}\in\ell^{2}$, $𝓡\langle\bm{u},\bm{\mathcal{R}}_{n}\rangle\to^{d}\langle\bm{u},\bm{\mathcal{R}}\rangle$ as $n\to\infty$.
@@ 3 Application: testing linear hypotheses
As an application of Theorem 2.3, we consider testing a linear hypothesis for the high-dimensional parameter $\bm{\theta}$. Let $q(\leq p)$ be an integer, $\bm{A}\in\mathbb{R}^{q\times p}$ a given matrix, $\bm{b}\in\mathbb{R}^{q}$ a given vector. Let $\bm{\Sigma}_{A}=\bm{A}_{[q],T_{\gamma}}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{A}_{[q],T_{\gamma}}^{\top}$, where $\bm{J}$ is defined in (2.1). Moreover, let $r$ be the rank of $\bm{\Sigma}_{A}$, and let $\Lambda_{1}\leq\ldots\leq\Lambda_{r}$ be the non-zero eigenvalues of $\bm{\Sigma}_{A}$. Since $s_{\gamma}$ is independent of $n$ and $p$, we have $r\leq\min(s_{\gamma},q)$, so that $r$ is fixed. To test $\mathcal{H}_{0}:\bm{A}\bm{\theta}=\bm{b}$ against $\mathcal{H}_{1}:\bm{A}\bm{\theta}\neq\bm{b}$, consider
@@
Remark 5.
An extension to the case $q=q_{n}\to\infty$ is also possible under an additional assumption that $\bm{A}$ defines a bounded linear operator on $\ell^{2}$.
@@
Let Assumptions 14 hold. Under $\mathcal{H}_{0}$,
$W_{n}\to^{d}\sigma^{2}\sum_{j=1}^{r}\Lambda_{j}\chi_{1,j}^{2}$
as $n\to\infty$, where $\chi_{1,1}^{2},\ldots,\chi_{1,r}^{2}$ are independent and $\chi^{2}(1)$-distributed random variables.
@@
$\hat{\bm{\Sigma}}_{nA}=\bm{A}_{[q],T_{\gamma}}\bm{J}_{nT_{\gamma},T_{\gamma}}^{-1}\bm{A}_{[q],T_{\gamma}}^{\top}.$
@@
Let Assumptions 14 hold. Then, $\hat{\sigma}_{n}^{2}\to^{p}\sigma_{0}^{2}$ and $\max_{j\in[r]}|\hat{\Lambda}_{n,j}-\Lambda_{j}|\to^{p}0$ as $n\to\infty$.
Let $c_{\alpha}$ and $\hat{c}_{\alpha}$ denote the upper $(1-\alpha)$-quantiles of the distributions of
@@
respectively; thus $\hat{c}_{\alpha}$ is a natural estimator of $c_{\alpha}$. We reject $\mathcal{H}_{0}$ if $W_{n}>\hat{c}_{\alpha}$, equivalently with test function $1_{\{W_{n}>\hat{c}_{\alpha}\}}$. The following theorem shows that this test attains asymptotic level $\alpha$ under $\mathcal{H}_{0}$.
@@
Let Assumptions 14 hold. Under $\mathcal{H}_{0}$, $\mathsf{P}(W_{n}>\hat{c}_{\alpha})\to\alpha$ as $n\to\infty$.
@@ 4 Proofs
Throughout this section, all limits refer to the regime $n\to\infty$ with $p=p_{n}$ possibly diverging.
@@
It suffices to show that $\hat{T}_{n}=T_{\gamma}$ under the event $\{\|\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{\infty}\leq c_{n}\}$. For every $j\in\hat{T}_{n}^{c}$, we have $|\hat{\theta}_{nj}|\leq\tau_{n}$ and $|\theta_{0j}|-|\hat{\theta}_{nj}|\leq|\hat{\theta}_{nj}-\theta_{0j}|\leq c_{n}$. Then, we have $|\theta_{0j}|\leq c_{n}+\tau_{n}\leq\gamma$, which implies that $j\in T_{\gamma}^{c}$. Therefore, we obtain $\hat{T}_{n}^{c}\subset T_{\gamma}^{c}$, or equivalently, $T_{\gamma}\subset\hat{T}_{n}$. It holds that $T_{\gamma}^{c}=T_{W}\cup T_{0}^{c}$. For every $j\in T_{0}^{c}$, we have $|\hat{\theta}_{nj}|=|\hat{\theta}_{nj}-\theta_{0j}|\leq c_{n}\leq\tau_{n}$, which implies that $j\in\hat{T}_{n}^{c}$. Similarly, for every $j\in T_{W}$, it holds that $|\hat{\theta}_{nj}|\leq c_{n}+|\theta_{0j}|\leq c_{n}+\eta_{p}\leq\tau_{n}$, which implies that $j\in\hat{T}_{n}^{c}$. In conclusion, we have $T_{\gamma}^{c}\subset\hat{T}_{n}^{c}$. ∎
@@
It follows from Bernstein’s inequality for sums of i.i.d. matrices established by Tropp (2012) that $\|\bm{J}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{T_{\gamma},T_{\gamma}}\|_{\mathrm{op}}=o_{p}(1)$. ∎
Proof of Lemma 2.4.
@@
$n^{1/2}\psi_{nT_{\gamma}}(\bm{\theta}_{0T_{\gamma}})1_{\{\hat{T}_{n}=T_{\gamma}\}}=\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{J}_{n\hat{T}_{n},\hat{T}_{n}}n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{\hat{T}_{n}}1_{\{\hat{T}_{n}=T_{\gamma}\}}.$
Since $\|\bm{J}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{T_{\gamma},T_{\gamma}}\|_{\mathrm{op}}=o_{p}(1)$ and $1_{\{\hat{T}_{n}=T_{\gamma}\}}=1+o_{p}(1)$, we obtain
$n^{1/2}\bm{u}_{T_{\gamma}}^{\top}\psi_{nT_{\gamma}}(\bm{\theta}_{0T_{\gamma}})1_{\{\hat{T}_{n}=T_{\gamma}\}}=\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{J}_{nT_{\gamma},T_{\gamma}}n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}1_{\{\hat{T}_{n}=T_{\gamma}\}}$
@@
Moreover,
$n^{-1}\sum_{i=1}^{n}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{iT_{\gamma}}\bm{Z}_{iT_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}}\epsilon_{i}^{2}\to^{p}\sigma^{2}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}}.$
Furthermore, since $\bm{Z}_{1},\ldots,\bm{Z}_{n}$ and $\epsilon_{1},\ldots,\epsilon_{n}$ are i.i.d. copies of sub-Gaussian variables $\bm{Z}$ and $\epsilon$, respectively, and $s_{\gamma}$ is independent of $n$ and $p$, we have $\mathsf{E}[|\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{T_{\gamma}}|^{2+\delta}]=O(1)$ and $\mathsf{E}[|\epsilon|^{2+\delta}]=O(1)$ for any $\delta>0$. Thus,
$\sum_{i=1}^{n}\,\mathsf{E}[|n^{-1/2}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{iT_{\gamma}}\epsilon_{i}|^{2+\delta}]\leq n^{-\delta/2}\,\mathsf{E}[|\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{T_{\gamma}}|^{2+\delta}]\,\mathsf{E}[|\epsilon|^{2+\delta}]\\ =o(1).$
@@
Let $\mathcal{H}_{0}$ hold. Define a linear operator $\mathcal{T}_{A}:\ell^{2}\to\ell^{2}$ by
$\langle\bm{e}_{j},\mathcal{T}_{A}(\bm{x})\rangle=\langle\bm{a}_{j},\bm{x}\rangle 1_{\{j\in[q]\}}\quad(j\in\mathbb{N})$
@@
First, we show the consistency of $\hat{\sigma}_{n}^{2}$. Decompose $\hat{\sigma}_{n}^{2}$ as $\hat{\sigma}_{n}^{2}=S_{n1}-2S_{n2}+S_{n3}$, where
@@
By the law of large numbers, $S_{n1}\to^{p}\sigma_{0}^{2}$. To prove $S_{n2}=o_{p}(1)$, it suffices to show that

### v2

@@ 2 Model set‑up and main result
For each $n\in\mathbb{N}$, we work with the triangular array $\{(Y_{i,n},\bm{Z}_{i,n})\}_{i\in[n]}$ satisfying
@@
where $\bm{\theta}_{n}\in\mathbb{R}^{p_{n}}$ and $\bm{Z}_{i,n}\in\mathbb{R}^{p_{n}}$ for $i\in[n]$. Assume that $\{(\bm{Z}_{i,n},\epsilon_{i,n})\}_{i\in[n]}$ are i.i.d. with $\mathsf{E}(\bm{Z}_{1,n})=\bm{0}$, $\mathsf{E}(\epsilon_{1,n})=0$, $\mathsf{V}(\epsilon_{1,n})=\sigma_{n}^{2}>0$, and $\bm{Z}_{1,n}$ independent of $\epsilon_{1,n}$. Let $\bm{\theta}_{0,n}$ denote the true parameter vector. Unless otherwise stated, throughout this paper all limits are taken as $n\to\infty$, allowing $p_{n}$ to diverge. Hereafter, we suppress the subscript $n$ when no confusion can arise. Write $\bm{\theta}_{0}=(\theta_{01},\ldots,\theta_{0p})^{\top}$. Define the index sets
@@
where $\theta_{\min}:=\min_{j\in T_{0}}|\theta_{0j}|$, $\gamma>0$ is a fixed constant independent of $n$ and $p$, and $\eta_{p}$ is a positive sequence depending on $p$. We write $s_{0}:=|T_{0}|$ for the support size, which may diverge, and $s_{\gamma}:=|T_{\gamma}|$ for the number of coefficients of moderate magnitude, kept fixed as $n$ varies.
@@
    $T_{\gamma}$ is fixed and independent of $n$ and $p$.
@@
Assumption 1-(ii) states that the non-zero coefficients outside $T_{\gamma}$ are precisely the weak signals, and Assumption 1-(iii) requires that their aggregate magnitude is asymptotically negligible. Under Assumption 1, $(\bm{\theta}_{0,n}^{\top},0,0,\ldots)\in\mathbb{R}^{\infty}$ lies in $\ell^{2}$ uniformly in $n$, because
$\|\bm{\theta}_{0,n}\|_{2}^{2}=\sum_{j\in T_{\gamma}}|\theta_{0j}|^{2}+\sum_{j\in T_{0}\setminus T_{\gamma}}|\theta_{0j}|^{2}\leq s_{\gamma}\theta_{\max}^{2}+s_{0}\eta_{p}^{2}=O(1).$
3
Let $\hat{\bm{\theta}}_{n}=(\hat{\theta}_{n1},\ldots,\hat{\theta}_{np})^{\top}$ be a sparse estimator for $\bm{\theta}$ such that $\mathsf{P}(\|\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{\infty}>r_{n})\to 0$ for some rate $r_{n}\to 0$. Given a threshold sequence $\{\tau_{n}\}_{n\in\mathbb{N}}$, define the sequence $\{\hat{T}_{n}\}_{n\in\mathbb{N}}$ of an estimator of $T_{0}$ by
@@
Under standard conditions, the lasso estimator and the Dantzig selector obey the high-probability bound $\|\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{\infty}\leq Cn^{-1/2}(\log{p})^{1/2}$ for some $C>0$. Accordingly, in such cases we may take $r_{n}\asymp n^{-1/2}(\log p)^{1/2}$.
@@
The following condition holds: $r_{n}+\eta_{p}<\tau_{n}<\gamma-r_{n}$.
@@
Let Assumptions 12 hold. Then, $\mathsf{P}(\hat{T}_{n}=T_{\gamma})\to 1$.
We define the ordinary least squares estimator $\tilde{\bm{\theta}}_{n}=(\tilde{\theta}_{n1},\ldots,\tilde{\theta}_{np})^{\top}$ after variable selection as the solution to the estimating equation
@@
$\hat{\bm{J}}_{n}:=n^{-1}\sum_{i=1}^{n}\bm{Z}_{i}\bm{Z}_{i}^{\top},\quad\bm{J}_{n}:=\,\mathsf{E}[\bm{Z}_{1}\bm{Z}_{1}^{\top}].$
We suppose that there exists a positive definite matrix $\bm{J}_{T_{\gamma},T_{\gamma}}=\lim\bm{J}_{nT_{\gamma},T_{\gamma}}$. In addition, we impose the following conditions.
@@
    There exists $\lambda_{*}$, independent of $n$ and $p$, such that $0<\lambda_{*}\leq\Lambda_{\min}(\bm{J}_{nT_{\gamma},T_{\gamma}})$.
@@
    $\sup_{n\in\mathbb{N}}\|\bm{Z}_{1,n}\|_{\psi_{2}}<\infty$.
@@
    $\sup_{n\in\mathbb{N}}\|\epsilon_{1,n}\|_{\psi_{2}}<\infty$.
@@
Let Assumptions 1 and 3 hold. Then, $\|\hat{\bm{J}}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{nT_{\gamma},T_{\gamma}}\|_{\mathrm{op}}=o_{p}(1)$.
@@
Fujimori and Tsukuda (2026) analyze asymptotic normality for high‑dimensional estimators via weak convergence in $\ell^{2}$, including settings with a possibly infinite‑dimensional nuisance parameter and dependent observations. Their analysis is confined to the fixed-sparsity regime, which materially simplifies the arguments. In contrast, we allow $s_{0}\to\infty$ while keeping $s_{\gamma}$ fixed.
@@ 3 Application: testing linear hypotheses
As a direct application of Theorem 2.3, we consider testing linear hypotheses for $\bm{\theta}$; the resulting test statistic and its null limit are stated below. Let $q(\leq p)$ be an integer, $\bm{A}\in\mathbb{R}^{q\times p}$ a given matrix, $\bm{b}\in\mathbb{R}^{q}$ a given vector. Let $\bm{\Sigma}_{A}=\bm{A}_{[q],T_{\gamma}}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{A}_{[q],T_{\gamma}}^{\top}$. Moreover, let $r$ be the rank of $\bm{\Sigma}_{A}$, and let $\Lambda_{1}\leq\ldots\leq\Lambda_{r}$ be the non-zero eigenvalues of $\bm{\Sigma}_{A}$. Since $s_{\gamma}$ is independent of $n$ and $p$, we have $r\leq\min(s_{\gamma},q)$, so that $r$ is fixed. To test $\mathrm{H}_{0}:\bm{A}\bm{\theta}=\bm{b}$ against $\mathrm{H}_{1}:\bm{A}\bm{\theta}\neq\bm{b}$, consider
@@
Let Assumptions 14 hold. Under $\mathrm{H}_{0}$,
$W_{n}\to^{d}\sigma^{2}\sum_{j=1}^{r}\Lambda_{j}\chi_{1,j}^{2},$
where $\chi_{1,1}^{2},\ldots,\chi_{1,r}^{2}$ are independent and $\chi^{2}(1)$-distributed random variables.
@@
$\hat{\bm{\Sigma}}_{nA}=\bm{A}_{[q],T_{\gamma}}\hat{\bm{J}}_{nT_{\gamma},T_{\gamma}}^{-1}\bm{A}_{[q],T_{\gamma}}^{\top}.$
@@
Let Assumptions 14 hold. Then, $\hat{\sigma}_{n}^{2}\to^{p}\sigma^{2}$ and $\max_{j\in[r]}|\hat{\Lambda}_{n,j}-\Lambda_{j}|\to^{p}0$.
Let $c_{\alpha}$ and $\hat{c}_{\alpha}$ denote the upper $\alpha$-quantiles of the distributions of
@@
respectively; thus $\hat{c}_{\alpha}$ is a natural estimator of $c_{\alpha}$. We reject $\mathrm{H}_{0}$ if $W_{n}>\hat{c}_{\alpha}$, equivalently with test function $\varphi_{n}=1_{\{W_{n}>\hat{c}_{\alpha}\}}$. The following theorem shows that this test attains asymptotic level $\alpha$ under $\mathrm{H}_{0}$.
@@
Let Assumptions 14 hold. Under $\mathrm{H}_{0}$, $\mathsf{P}(\varphi_{n}=1)\to\alpha$.
@@
Remark 5.
Zhu and Bradic (2018) and Zhang et al. (2025) conduct statistical inference with dense $\bm{\theta}$ by imposing sparsity on orthogonalized loading directions obtained after suitable projection. By contrast, we impose no sparsity on $\bm{A}$ and instead proceed under stronger assumptions on $\bm{\theta}$, thereby situating our results in a different inferential setting.
Remark 6.
Although we assume $q$ is fixed, the results can be extended to the regime $q=q_{n}\to\infty$,provided that $\bm{A}$ defines a bounded linear operator on $\ell^{2}$.
@@ 4 Proofs
It suffices to show that $\hat{T}_{n}=T_{\gamma}$ on the event $\{\|\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{\infty}\leq r_{n}\}$. For every $j\in\hat{T}_{n}^{c}$, we have $|\hat{\theta}_{nj}|\leq\tau_{n}$ and $|\theta_{0j}|-|\hat{\theta}_{nj}|\leq|\hat{\theta}_{nj}-\theta_{0j}|\leq r_{n}$. Then, we have $|\theta_{0j}|\leq r_{n}+\tau_{n}\leq\gamma$, which implies that $j\in T_{\gamma}^{c}$. Therefore, we obtain $\hat{T}_{n}^{c}\subset T_{\gamma}^{c}$, or equivalently, $T_{\gamma}\subset\hat{T}_{n}$. It holds that $T_{\gamma}^{c}=T_{W}\cup T_{0}^{c}$. For every $j\in T_{0}^{c}$, we have $|\hat{\theta}_{nj}|=|\hat{\theta}_{nj}-\theta_{0j}|\leq r_{n}\leq\tau_{n}$, which implies that $j\in\hat{T}_{n}^{c}$. Similarly, for every $j\in T_{W}$, it holds that $|\hat{\theta}_{nj}|\leq r_{n}+|\theta_{0j}|\leq r_{n}+\eta_{p}\leq\tau_{n}$, which implies that $j\in\hat{T}_{n}^{c}$. In conclusion, we have $T_{\gamma}^{c}\subset\hat{T}_{n}^{c}$. ∎
@@
The conclusion follows from the law of large numbers for triangular arrays. ∎
Proof of Theorem 2.3.
It suffices to establish the following (i) and (ii); see, for example, Chapter 1.8 of van der Vaart and Wellner (2023): (i) For every $\zeta>0$, $\mathsf{P}(\|n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}^{c}}\|_{2}^{2}>\zeta)\to 0$. (ii) For every $\bm{u}\in\ell^{2}$, $𝓡\langle\bm{u},\bm{\mathcal{R}}_{n}\rangle\to^{d}\langle\bm{u},\bm{\mathcal{R}}\rangle$.
@@
$n^{1/2}\psi_{nT_{\gamma}}(\bm{\theta}_{0T_{\gamma}})1_{\{\hat{T}_{n}=T_{\gamma}\}}=\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\hat{\bm{J}}_{n\hat{T}_{n},\hat{T}_{n}}n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{\hat{T}_{n}}1_{\{\hat{T}_{n}=T_{\gamma}\}}.$
Since $\|\hat{\bm{J}}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{T_{\gamma},T_{\gamma}}\|_{\mathrm{op}}\leq\|\hat{\bm{J}}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{nT_{\gamma},T_{\gamma}}\|_{\mathrm{op}}+\|\bm{J}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{T_{\gamma},T_{\gamma}}\|_{\mathrm{op}}=o_{p}(1)$ and $1_{\{\hat{T}_{n}=T_{\gamma}\}}=1+o_{p}(1)$, we obtain
$n^{1/2}\bm{u}_{T_{\gamma}}^{\top}\psi_{nT_{\gamma}}(\bm{\theta}_{0T_{\gamma}})1_{\{\hat{T}_{n}=T_{\gamma}\}}=\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\hat{\bm{J}}_{nT_{\gamma},T_{\gamma}}n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}1_{\{\hat{T}_{n}=T_{\gamma}\}}$
@@
Moreover, the sub-Gaussian bound implies
$n^{-1}\sum_{i=1}^{n}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{iT_{\gamma}}\bm{Z}_{iT_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}}\epsilon_{i}^{2}\to^{p}\sigma^{2}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}},$
because
$n^{-1}\sum_{i=1}^{n}\,\mathsf{E}[\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{iT_{\gamma}}\bm{Z}_{iT_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}}\epsilon_{i}^{2}]=\sigma^{2}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{J}_{nT_{\gamma},T_{\gamma}}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}}\to\sigma^{2}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}}.$
Furthermore, since $s_{\gamma}$ is independent of $n$ and $p$, the sub-Gaussian bound implies $\mathsf{E}[|\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{1T_{\gamma}}|^{2+\delta}]=O(1)$ and $\mathsf{E}[|\epsilon_{1}|^{2+\delta}]=O(1)$ for any $\delta>0$. Thus,
$\sum_{i=1}^{n}\,\mathsf{E}[|n^{-1/2}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{iT_{\gamma}}\epsilon_{i}|^{2+\delta}]\leq n^{-\delta/2}\,\mathsf{E}[|\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{1T_{\gamma}}|^{2+\delta}]\,\mathsf{E}[|\epsilon_{1}|^{2+\delta}]\\ =o(1).$
@@
Let $\mathrm{H}_{0}$ hold. Define a linear operator $\mathcal{T}_{A}:\ell^{2}\to\ell^{2}$ by
$\langle\bm{e}_{j},\mathcal{T}_{A}(\bm{x})\rangle=\begin{cases}\langle\bm{a}_{j},\bm{x}\rangle&(j\in[q])\\ 0&(j\in\mathbb{N}\setminus[q])\end{cases}$
@@
First, we show $\hat{\sigma}_{n}^{2}\to^{p}\sigma^{2}$. Decompose $\hat{\sigma}_{n}^{2}$ as $\hat{\sigma}_{n}^{2}=S_{n1}-2S_{n2}+S_{n3}$, where
@@
By the law of large numbers, $S_{n1}\to^{p}\sigma^{2}$. To prove $S_{n2}=o_{p}(1)$, it suffices to show that